The Right Dipole was increasingly unstable. There appeared to be some correlation with readback issues with the NMR. I thought it might be a problem with the software regulator, so I rebooted the IOC. This did not help. The current was wandering upwards. I went into the hall to look with the scope at what I could. The supply still was not responding. So, I hit the "big red button" and dumped the magnet. Looking in the Archiver there is some indication the instability began around 16:30 today and the current started increasing around 19:15. Reboot was around 19:26. The current was not making a clean ramp, just wandering around. It would seem the regulator circuit is having some issue. As a wild guess I would say 4 to 6 hours to trouble shoot (three hours to convince Javier to lay hands on it while Howard recites the incantation, one to two hours to perform the rite). But, seriously, 4 to 6 hours is just my rough guess. Till we can go in to work on it during a day shift, run with the Left Arm only. Also, the background was around 7mR around the supply. Radcon will be required to set the guidelines to do the work.